# Hypothesis-Driven Theory-of-Mind Reasoning for Large Language Models

### v1

@@ Abstract
Existing LLM reasoning methods have shown impressive capabilities across various tasks, such as solving math and coding problems. However, applying these methods to scenarios without ground-truth answers or rule-based verification methods—such as tracking the mental states of an agent—remains challenging. Inspired by the sequential Monte Carlo algorithm, we introduce thought-tracing, an inference-time reasoning algorithm designed to trace the mental states of specific agents by generating hypotheses and weighting them based on observations without relying on ground-truth solutions to questions in datasets. Our algorithm is modeled after the Bayesian theory-of-mind framework, using LLMs to approximate probabilistic inference over agents’ evolving mental states based on their perceptions and actions. We evaluate thought-tracing on diverse theory-of-mind benchmarks, demonstrating significant performance improvements compared to baseline LLMs. Our experiments also reveal interesting behaviors of the recent reasoning models – e.g., o1 and R1 – on theory-of-mind, highlighting the difference of social reasoning compared to other domains.
@@ 1 Introduction
New reasoning algorithms and training paradigms for large language models (LLMs) have recently achieved remarkable breakthroughs in domains where well-defined ground-truth answers are readily available, enabling partial-solution evaluation and objective verification – e.g., mathematics, coding, and puzzles (Yao et al., 2023; Besta et al., 2024; OpenAI et al., 2024b; DeepSeek-AI et al., 2025b). This stands in stark contrast to social reasoning, where objective answers are not easily obtainable, making partial-solution evaluation less straightforward. Furthermore, the information-asymmetric nature of theory-of-mind (ToM) – which requires inferring hidden mental states – makes social reasoning more challenging due to its increased uncertainty (Zhi-Xuan et al., 2024b; Nafar et al., 2024).
Figure 1: An illustrative overview of our thought-tracing algorithm. The original context is parsed into a trajectory of the target agent. At each time step, the algorithm generates multiple hypotheses about the agent’s beliefs and weighs them based on the likelihood of the agent’s action. These weights are then used to resample the hypotheses, prioritizing more promising ones to be used for the next time step. Further details are in Section 2.
@@
To this end, we propose thought-tracing, an inference-time reasoning algorithm for LLMs that can infer and track the mental states of target agents. Our method is modeled after the Bayesian theory-of-mind framework (BToM; Baker et al., 2017), viewing the input context as a sequence of states and agent actions, and treating ToM reasoning as probabilistic inference about the agent’s mental states based on their perceptions and actions. To account for the inherent uncertainty in this task, we follow the high-level structure of sequential Monte Carlo inference (SMC; Del Moral et al., 2006; Lew et al., 2023b), tracking multiple weighted hypotheses about the agent’s mental states. Importantly, however, these hypotheses are represented in open-ended natural language, and are both generated and weighted by LLMs. By using LLMs in this way, our method offers a generalizable approach to uncovering the underlying thoughts that drive agent behavior, without requiring the explicit probabilistic models assumed by BToM algorithms, and without relying on question-answer annotations or benchmark-specific assumptions.
Although thought-tracing is not specifically designed for solving benchmark questions, we evaluate the algorithm using state-of-the-art LLMs on four theory-of-mind benchmarks to assess whether the traced thoughts can enhance downstream question answering task performance. Experiments show that (1) thought-tracing consistently improves performance on all tested models across all benchmarks, (2) additional inference-time compute (e.g., chain-of-thought) further aids models to process contexts interleaved with mental states, and (3) models with thought-tracing outperform reasoning models (e.g., o1 and R1) despite using significantly shorter reasoning traces.
@@
These results suggest that thought-tracing represents a promising step towards more robust inference-time ToM reasoning. We aim to spark new discussions on inference-time reasoning in the social domain, contrasting with the predominant focus on math and coding.
2 Thought-tracing
We introduce thought-tracing, an inference-time algorithm that performs ToM reasoning by inferring and propagating weighted hypotheses about agents’ thoughts. These hypotheses are represented in natural language, and are generated and weighted using LLMs. Thought-tracing follows the conceptual structure of Bayesian Theory of Mind, and takes inspiration from sequential Monte Carlo to perform approximate inference over agents’ mental states.
@@ 2.1 Background
Thought-tracing follows the structure of this framework without using explicit models of rational belief updating and action selection. Instead, we use LLMs to simulate how an agent is likely to update their mental states in response to what they perceive, and to evaluate the likelihood of an agent’s actions given their mental states. This enables greater generality, and decomposes social inference into the simpler tasks of forward simulation and likelihood evaluation by LLMs.
Sequential Monte Carlo (SMC) refers to a family of algorithms designed for incremental inference over sequences of posterior distributions (Del Moral et al., 2006), such as posterior inference over latent dynamics from time series data. SMC uses a collection of weighted hypotheses (called particles) to approximate each distribution in the sequence. Given particles for step $t-1$ in the sequence, SMC generates particles for step $t$ via propagation (extending each previous particle with new latent states that exist at step $t$) and reweighting (weighting each particle by the likelihood of the observed data under that particle’s latent states). The particles are then resampled according to their weights (focusing samples on regions of high posterior probability) and optionally rejuvenated with Markov chain Monte Carlo (MCMC) to increase particle diversity (Chopin, 2002). To improve inference quality, SMC can make use of custom proposals at for propagation or rejuvenation (Lew et al., 2023a), using data-driven cues to generate more plausible hypotheses (Perov et al., 2015).
To infer hidden mental states that change over time, thought-tracing follows an SMC-like structure, using LLMs as proposals for propagating hypotheses about agents’ mental states, and weighting these hypotheses by likelihood scores generated by LLMs. However, for simplicity, thought-tracing does not compute full importance weights for each particle, since this would require accessing LLM log probabilities not provided by most APIs.
@@ 2.2 Algorithm
16:     return $\tilde{\tau}$triangleright $\tau$ with traced thoughts
Algorithm 1 shows the pseudo-code for thought-tracing. Specifically, it leverages sequential Monte Carlo principles to maintain a dynamically updated set of natural language hypotheses about the target agent’s mental state, weighting them in response to new observations. These hypotheses are updated based on the agent’s perception, prior beliefs, and actions.
@@
Thought-tracing begins by sampling a collection of weighted hypotheses $H_{1}=\{(h_{1}^{(i)},w_{1}^{(i)})\}_{i=1}^{N}$, where $h_{1}^{(i)}$ denotes a hypothesis and $w_{1}^{(i)}$ is the associated weight, from an initial distribution that approximates the step-$1$ posterior $p(h_{1}|s_{1},a_{1},p_{1})$. In practice, the LLM is given $(s_{1},a_{1},p_{1})$ as input, and produces a list of $N$ hypotheses, with each hypothesis assigned an initial weight of $w_{1}^{(i)}=\frac{1}{N}$. We set $N=4$ for our experiments.
@@
At every step $t$, each hypothesis $h_{t-1}^{(i)}$ is propagated forward to produce a new hypothesis $h_{t}^{(i)}$ based on the trajectory so far:
$h_{t}^{(i)}\sim p_{\theta}\bigl{(}h_{t}\,\big{|}\;h_{t-1}^{(i)},\{(s_{\tau},a_{\tau},p_{\tau})\}_{\tau=1}^{t}\bigr{)}.$ (1)
Here, $\theta$ denotes the parameters of an LLM and $\{(s_{j},a_{j},p_{j})\}_{j=1}^{t}$ denotes the trajectory (i.e., state, action, perception) up to time $t$. In effect, the LLM is used as a data-driven proposal that generates a plausible mental state $h_{t}$ given the observations.
@@
After propagation, each hypothesis $h_{t}^{(i)}$ receives a weight update based on how well it explains the action or utterance $a_{t}$. The LLM is used to output the likelihood that $a_{t}$ would be produced under each hypothesis:
$w_{t}^{(i)}=p_{\theta}\bigl{(}a_{t}\big{|}h_{t}^{(i)},\{(s_{j},a_{j},p_{j})\}_{j=1}^{t-1},s_{t},p_{t}\bigr{)}$ (2)
While it is possible to compute this likelihood from model’s logprobs (as done in Zhi-Xuan et al. (2024b)), in practice we find that instructing the LLM to choose from six options—from “very likely (around 90%)” to “very unlikely (below 10%)”—yields better performance and stability. We map these options to numerical scores and then normalize the weights so that $\sum_{i}^{N}w_{t}^{(i)}=1$.
@@
The final output of thought-tracing is a trajectory of the target agent $\mathcal{A}$ with the summarized hypotheses interleaved. An example is in Table 1. State, action, perception, and beliefs of $\mathcal{A}$ at each time step is concatenated as a reasoning trace formatted with HTML-style tags.
An Example Traced Thought
<context 2> He then proceeds to open the third Room C cabinet and closes it as well. <note>David perceived that the third Room C cabinet is empty.</note> <David’s updated thoughts>David believed that the third Room C cabinet might contain an Object E. This belief is inferred from his action of opening the cabinet to check its contents, suggesting that he thought there was a possibility of finding Object E there. However, upon opening it, he perceived that it was empty, confirming that his belief was incorrect regarding the presence of an Object E in that cabinet.</David’s updated thoughts> </context 2>
<context 3> Following this, he opens and closes the stove. <note>David perceived the stove in Room C, which contained an Object H and an Object J, as he opened and then closed it.</note> <David’s updated thoughts>David believed that there might be an Object E in the stove. This belief is inferred from his systematic search through the storage spaces in Room C, as he had not yet found the Object E he was looking for in the cabinets. His actions of opening and closing the stove suggest that he was checking to see if it contained the desired object. However, upon opening it, he perceived that it contained an Object H and a piece of Object J, confirming that his belief regarding the presence of an Object E in the stove was incorrect.</David’s updated thoughts> </context 3>
Table 1: An example traced thoughts from our thought-tracing algorithm on MMToM-QA.
@@ 3 Experiments
Our thought-tracing algorithm offers a general method for tracing the mental states of a target agent, given context. While the algorithm is not explicitly designed to solve benchmark questions, we evaluate its validity by testing whether it can improve existing models’ performance on theory-of-mind benchmarks. We apply thought-tracing to off-the-shelf LLMs and feed the generated thought traces along with the original context and compare their performance against other baselines, such as reasoning models.
Benchmarks & Metrics
We test our approach on four theory-of-mind benchmarks that span diverse format and question types:
@@
(2) BigToM (Gandhi et al., 2023) is a benchmark featuring synthetic narratives centered on a person’s desires, actions, and beliefs, along with an event that changes the state of the environment. It includes questions about the person’s beliefs and actions. We find some of the narratives and ground-truth answers in BigToM include incoherent stories and errors. To address this issue, we randomly sampled 600 scenarios from the dataset and re-annotated by crowdsourcing 510 online participants (See details in Appendix A). We only retained samples with an agreement rate higher than 90%, which resulted in 249 samples. We report the average accuracy of the questions.
(3) FANToM (Kim et al., 2023) is a multi-party conversation question-answering dataset designed to test coherent theory-of-mind capabilities. In FANToM, the speakers join and leave the conversation while it continues, making the participants hold both false and true beliefs. The benchmark includes first-order and second-order theory-of-mind questions about the beliefs of conversation participants in various formats, such as multiple-choice and list type questions. We use 64 sampled conversations from the short version of FANToM, containing a total of 1,086 questions. We report three metrics (i) All Qs, (ii) Answerability All Qs, and (iii) Info-Accessibility All Qs, each requiring the model to correctly answer all questions for the corresponding question types for a given conversation snippet.
@@
Table 2: Results across four theory-of-mind benchmarks. The underlined metric indicates the primary metric for each benchmark. The notation ‘+TT’ denotes that our thought-tracing method has been applied, while ‘+CoT’ indicates that chain-of-thought has been used.
@@
We apply our method to four state-of-the-art LLMs and compare their performance, including five reasoning models: GPT-4o (OpenAI et al., 2024a), o1, o1-mini (OpenAI et al., 2024b), o3-mini (OpenAI, 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025a), Llama 3.3 70B Instruct (Dubey et al., 2024), Gemini 1.5 Pro (Gemini-Team et al., 2024), Qwen 2.5 72B (Yang et al., 2024), and QwQ 32B preview (Qwen-Team, 2024). We also compare with zero-shot chain-of-thought (CoT) reasoning (Kojima et al., 2022).
@@ 3.1 Overall Results
Thought-tracing (+TT) improves all baseline models across all benchmarks.
Table 2 shows the results on all four benchmarks. Although thought-tracing does not take benchmark questions as input, its output significantly improves the model performance on the main metrics – the average accuracy (Avg.) on Paraphrased ToMi, BigToM, MMToM-QA, and the AllQs score on FANToM. For example, Llama 3.3 70B + TT, Gemini 1.5 Pro + TT, and Qwen 2.5 72B + TT significantly outperforms GPT-4o + CoT on Paraphrased ToMi. These results indicate that the traced mental states from thought-tracing provide accurate intermediate reasoning steps that helps models correctly answer mental state related questions.
Thought-tracing guides better reasoning trajectories.
The fact that thought-tracing outperforms CoT indicates the traced thoughts exhibit better reasoning traces. Moreover, when CoT is applied on top of the traced thoughts, models’ performance improves further in many cases. For example, Llama 3.3 70B and Qwen 2.5 72B achieve the first and second best scores on Paraphrased ToMi, respectively. All models achieves the best score when thought-tracing and CoT is applied sequentially on MMToM-QA. Interestingly, Qwen 2.5 does not show any improvement on FANToM scoring near 0, when thought-tracing is applied. However, when CoT is further applied its performance jumps to 32.1, even outperforming o1-medium-effort and o3-mini. These results indicate that results from thought-tracing can further guide the model to obtain better reasoning trajectories.

### v2

@@ Abstract
Existing LLM reasoning methods have shown impressive capabilities across various tasks, such as solving math and coding problems. However, applying these methods to scenarios without ground-truth answers or rule-based verification methods—such as tracking the mental states of an agent—remains challenging. Inspired by the sequential Monte Carlo algorithm, we introduce ThoughtTracing, an inference-time reasoning algorithm designed to trace the mental states of specific agents by generating hypotheses and weighting them based on observations without relying on ground-truth solutions to questions in datasets. Our algorithm is inspired by the Bayesian theory-of-mind framework, using LLMs to approximate probabilistic inference over agents’ evolving mental states based on their perceptions and actions. We evaluate ThoughtTracing on diverse theory-of-mind benchmarks, demonstrating significant performance improvements compared to baseline LLMs. [^1] Data and code are available at https://hyunw.kim/thought-tracing Our experiments also reveal interesting behaviors of the recent reasoning models – e.g., o3 and R1 – on theory-of-mind, highlighting the difference of social reasoning compared to other domains.
@@ 1 Introduction
New reasoning algorithms and training paradigms for large language models (LLMs) have recently achieved remarkable breakthroughs in domains where well-defined ground-truth answers are readily available, enabling partial-solution evaluation and objective verification – e.g., mathematics, coding, and puzzles (Yao et al., 2023; Besta et al., 2024; OpenAI, 2024b; DeepSeek-AI, 2025b). This stands in stark contrast to social reasoning, where objective answers are not easily obtainable, making partial-solution evaluation less straightforward. Furthermore, the information-asymmetric nature of theory-of-mind (ToM) – which requires inferring hidden mental states – makes social reasoning more challenging due to its increased uncertainty (Zhi-Xuan et al., 2024b; Nafar et al., 2024).
Figure 1: An illustrative overview of our ThoughtTracing algorithm. The original context is parsed into a trajectory of the target agent. At each time step, the algorithm generates multiple hypotheses about the agent’s beliefs and weighs them based on the likelihood of the agent’s action. These weights are then used to resample the hypotheses, prioritizing more promising ones to be used for the next time step. Further details are in Section 2.
@@
To this end, we propose ThoughtTracing, an inference-time reasoning algorithm for LLMs that can infer and track the mental states of target agents. Our method is conceptually inspired by the Bayesian theory-of-mind framework (BToM; Baker et al., 2017), viewing the input context as a sequence of states and agent actions, and treating ToM reasoning as probabilistic inference about the agent’s mental states based on their perceptions and actions. To account for the inherent uncertainty in this task, we follow the high-level structure of sequential Monte Carlo inference (SMC; Del Moral et al., 2006; Lew et al., 2023b), tracking multiple weighted hypotheses about the agent’s mental states. Importantly, however, these hypotheses are represented in open-ended natural language, and are both generated and weighted by LLMs. By using LLMs in this way, our method offers a generalizable approach to uncovering the underlying thoughts that drive agent behavior, without requiring the explicit probabilistic models assumed by BToM algorithms, and without relying on question-answer annotations or benchmark-specific assumptions.
Although ThoughtTracing is not specifically designed for solving benchmark questions, we evaluate the algorithm using state-of-the-art LLMs on four theory-of-mind benchmarks to assess whether the traced thoughts can enhance downstream question answering task performance. Experiments show that (1) ThoughtTracing consistently improves performance on all tested models across all benchmarks, (2) additional inference-time compute (e.g., chain-of-thought) further aids models to process contexts interleaved with mental states, and (3) models with ThoughtTracing outperform reasoning models (e.g., o3-mini and R1) despite using significantly shorter reasoning traces.
@@
These results suggest that ThoughtTracing represents a promising step towards more robust inference-time ToM reasoning. We aim to spark new discussions on inference-time reasoning in the social domain, contrasting with the predominant focus on math and coding.
2 ThoughtTracing
We introduce ThoughtTracing, an inference-time algorithm that performs ToM reasoning by inferring and propagating weighted hypotheses about agents’ thoughts. These hypotheses are represented in natural language, and are generated and weighted using LLMs. ThoughtTracing follows the conceptual structure of Bayesian Theory of Mind, and takes inspiration from sequential Monte Carlo to perform approximate inference over agents’ mental states.
@@ 2.1 Background
ThoughtTracing follows this framework’s structure without using explicit models of rational belief updating and action selection. Instead, we use LLMs to simulate how an agent is likely to update their mental states in response to their perceptions, and to evaluate the likelihood of an agent’s actions given their mental states. This enables greater generality, decomposing social inference into simpler tasks: forward simulation and likelihood evaluation by LLMs.
Sequential Monte Carlo (SMC) refers to a family of algorithms designed for incremental inference over sequences of posterior distributions (Del Moral et al., 2006), such as posterior inference over latent dynamics from time series data. SMC uses a collection of weighted hypotheses (called particles) to approximate each distribution in the sequence. Given particles for step t1t-1 in the sequence, SMC generates particles for step $t$ via propagation (extending each previous particle with new latent states that exist at step $t$) and reweighting (weighting each particle by the likelihood of the observed data under that particle’s latent states). The particles are then resampled according to their weights (focusing samples on regions of high posterior probability) and optionally rejuvenated with Markov chain Monte Carlo (MCMC) to increase particle diversity (Chopin, 2002). To improve inference quality, SMC can make use of custom proposals at for propagation or rejuvenation (Lew et al., 2023a), using data-driven cues to generate more plausible hypotheses (Perov et al., 2015).
To infer hidden mental states that change over time, ThoughtTracing follows an SMC-like structure, using LLMs as proposals for propagating hypotheses about agents’ mental states, and weighting these hypotheses by likelihood scores generated by LLMs. However, for simplicity, ThoughtTracing does not compute full importance weights for each particle, since this would require accessing LLM log probabilities not provided by most APIs.
@@ 2.2 Algorithm
16:  return $\tilde{\tau} \triangleright$ $\tau$ with traced thoughts
ThoughtTracing leverages sequential Monte Carlo principles to maintain a dynamically updated set of natural language hypotheses about the target agent’s mental state, weighting them in response to new observations. These hypotheses are updated based on the agent’s perception, prior beliefs, and actions. See Algorithm 1.
@@
ThoughtTracing begins by sampling a collection of weighted hypotheses $H_{1}=\{(h_{1}^{(i)},w_{1}^{(i)})\}_{i=1}^{N}$, where $h_{1}^{(i)}$ denotes a hypothesis and $w_{1}^{(i)}$ is the associated weight, from an initial distribution that approximates the step-$1$ posterior $p(h_{1}|s_{1},a_{1},p_{1})$. In practice, the LLM is given $(s_{1},a_{1},p_{1})$ as input, and produces a list of $N$ hypotheses, with each hypothesis assigned an initial weight of $w_{1}^{(i)}=\frac{1}{N}$. We set $N=4$ for our experiments.
@@
At every step $t$, each hypothesis $h_{t-1}^{(i)}$ is propagated forward to produce a new hypothesis $h_{t}^{(i)}$ based on the trajectory so far: $h_{t}^{(i)}\sim p_{\theta}\bigl{(}h_{t}\,\big{|}\;h_{t-1}^{(i)},\{(s_{\tau},a_{\tau},p_{\tau})\}_{\tau=1}^{t}\bigr{)}$. Here, $\theta$ denotes the parameters of an LLM and $\{(s_{j},a_{j},p_{j})\}_{j=1}^{t}$ denotes the trajectory (i.e., state, action, perception) up to time $t$. In effect, the LLM is used as a data-driven proposal that generates a plausible mental state $h_{t}$ given the observations.
@@
After propagation, each hypothesis $h_{t}^{(i)}$ receives a weight update based on how well it explains the action or utterance $a_{t}$. The LLM is used to output the likelihood that $a_{t}$ would be produced under each hypothesis: $w_{t}^{(i)}:=p_{\theta}\bigl{(}a_{t}\big{|}h_{t}^{(i)},\{(s_{j},a_{j},p_{j})\}_{j=1}^{t-1},s_{t},p_{t}\bigr{)}$. While it is possible to compute this likelihood from a model’s log probabilities (as done in Zhi-Xuan et al. (2024b)), in practice we find that instructing the LLM to choose from six options—from “very likely (around 90%)” to “very unlikely (below 10%)”—yields better performance and stability. Moreover, most closed-source APIs (e.g., GPT-4, Gemini) do not provide access to token-level log probabilities. We map these options to numerical scores and normalize so that $\sum_{i}^{N}\!w_{t}^{(i)}\!=\!1$.
@@
The final output of ThoughtTracing is a trajectory of the target agent $\mathcal{A}$ with the summarized hypotheses interleaved. An example is in Table 1. State, action, perception, and beliefs of $\mathcal{A}$ at each time step is concatenated as a reasoning trace formatted with HTML-style tags.
The full prompts used in ThoughtTracing can be found in the Appendix C. A case study on how the hypotheses and weights evolve over time is in the Appendix D.
<context 2>
He then proceeds to open the third Room C cabinet and closes it as well.
<note>David perceived that the third Room C cabinet is empty.</note>
<David’s updated thoughts>David believed that the third Room C cabinet might contain an Object E. This belief is inferred from his action of opening the cabinet to check its contents, suggesting that he thought there was a possibility of finding Object E there. However, upon opening it, he perceived that it was empty, confirming that his belief was incorrect regarding the presence of an Object E in that cabinet.</David’s updated thoughts>
</context 2>
<context 3>
Following this, he opens and closes the stove.
<note>David perceived the stove in Room C, which contained an Object H and an Object J, as he opened and then closed it.</note>
<David’s updated thoughts>David believed there might be an Object E in the stove. This belief is inferred from his systematic search through the storage spaces in Room C, as he had not yet found the Object E he was looking for in the cabinets. His actions of opening and closing the stove suggest that he was checking to see if it contained the desired object. However, upon opening it, he perceived that it contained an Object H and a piece of Object J, confirming that his belief regarding the presence of an Object E in the stove was incorrect.</David’s updated thoughts>
</context 3>
Table 1: An example traced thoughts from our ThoughtTracing algorithm on MMToM-QA.
@@ 3 Experiments
Our ThoughtTracing algorithm offers a general method for tracing the mental states of a target agent, given context. While the algorithm is not explicitly designed to solve benchmark questions, we evaluate its validity by testing whether it can improve existing models’ performance on four theory-of-mind (ToM) benchmarks. We apply ThoughtTracing to off-the-shelf LLMs and feed the generated thought traces along with the original context and compare their performance against other baselines, such as reasoning models.
@@
(2) BigToM (Gandhi et al., 2023) is a benchmark featuring synthetic narratives centered on a person’s desires, actions, and beliefs, along with an event that changes the environment’s state. It includes questions about the person’s beliefs and actions. We find some of the narratives and ground-truth answers in BigToM include incoherent stories and errors. To address this issue, we randomly sampled 600 scenarios from the dataset and re-annotated by crowdsourcing 510 online participants (see App. A). We only retained samples with an agreement rate higher than 90% (249 samples). We report the average accuracy of the questions.
(3) FANToM (Kim et al., 2023) is a multi-party conversation question-answering dataset designed to test coherent theory-of-mind capabilities. In FANToM, the speakers join and leave the conversation while it continues, making the participants hold both false and true beliefs. The benchmark includes first-order and second-order theory-of-mind questions about the beliefs of conversation participants in various formats, such as multiple-choice and list type questions. We use 64 sampled conversations from the short version of FANToM, containing a total of 1,086 questions. We report three metrics: (i) All Qs, (ii) Answerability All Qs, and (iii) Info-Accessibility All Qs, each requiring the model to correctly answer all questions for the corresponding question types for a given conversation snippet.
@@
Table 2: Results across four theory-of-mind benchmarks. The underlined metric indicates the primary metric for each benchmark. The notation ‘+$T$’ denotes that our ThoughtTracing method has been applied, while ‘+CoT’ indicates that chain-of-thought has been used. Full evaluation results can be found in Table 6 in Appendix.
@@
We apply our method to four state-of-the-art LLMs and compare their performance, and also compare with five recent reasoning models: GPT-4o (OpenAI, 2024a), o1, o1-mini (OpenAI, 2024b), o3-mini (OpenAI, 2025), DeepSeek-R1 (DeepSeek-AI, 2025a), Llama 3.3 70B Instruct (Dubey et al., 2024), Gemini 1.5 Pro (Gemini-Team, 2024), Qwen 2.5 72B (Yang et al., 2024), and QwQ 32B preview (Qwen-Team, 2024). We also compare with zero-shot chain-of-thought (CoT) reasoning (Kojima et al., 2022).
@@ 3.1 Overall Results
ThoughtTracing (+TT) improves all baseline models across all benchmarks.
Table 2 shows the results on all four benchmarks. Although ThoughtTracing does not take benchmark questions as input, its output significantly improves the model performance on the main metrics – the average accuracy (Avg.) on ParaphrasedToMi, BigToM, MMToM-QA, and the AllQs score on FANToM. For example, Llama 3.3 70B + TT, Gemini 1.5 Pro + TT, and Qwen 2.5 + TT significantly outperforms GPT-4o + CoT on ParaphrasedToMi. These results indicate that the traced mental states from ThoughtTracing provide accurate intermediate reasoning steps that helps models correctly answer mental state related questions.
ThoughtTracing guides better reasoning trajectories.
The fact that ThoughtTracing outperforms CoT indicates the traced thoughts exhibit better reasoning traces. Moreover, when CoT is applied on top of the traced thoughts, models’ performance improves further in many cases. For example, Llama 3.3 and Qwen 2.5 achieve the first and second best scores on ParaphrasedToMi, respectively. All models achieve the best score when ThoughtTracing and CoT is applied sequentially on MMToM-QA. Interestingly, Qwen 2.5 does not show any improvement on FANToM, when ThoughtTracing is applied. However, when CoT is further applied its performance jumps to 32.1, even outperforming o1-medium-effort and o3-mini.
On one hand, we find that, for GPT-4o and Gemini 1.5 Pro, the combination of +CoT+TT underperforms compared to +TT alone on ParaphrasedToMi. This underperformance is linked to the degradation on True Belief scenarios. In these scenarios, there is no information asymmetry—i.e., all participants share the same belief aligned with the ground-truth—making over-reasoning potentially harmful. Notably, vanilla CoT alone significantly reduces their performance on True Belief questions, suggesting that the extra reasoning steps may introduce noise or unnecessary complexity. This effect appears to be significantly stronger in closed-weight models (e.g., GPT-4o, Gemini 1.5 Pro) than open-weight ones (e.g., Llama 3.3, Qwen 2.5). Interestingly, the degradation of TT + CoT relative to TT alone is only observed in these closed-weight models, hinting at possible post-training differences.